OpenStudy (anonymous):

can some one show me the steps to this in sq rt . 2x^2-2x-5=0

OpenStudy (anonymous):

can anyone help me and show me the steps to this in sq rt . 2x^2-2x-5=0

OpenStudy (anonymous):

Quadratic formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Quadratic formula 2x^2-2x-5=0

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Quadratic_equation that is a quadratic equation! check out the website

OpenStudy (anonymous):

a=2,b=-2,c=-5. delta=b^2-4ac=44. x=(-b±√44)/2a=(1±√11)/2

OpenStudy (anonymous):

thank you